OpenStudy (anonymous):

Find the value of cos11π /6

OpenStudy (anonymous):

you have it

OpenStudy (anonymous):

its a + though not ÷ sin^2 + cos^2 = 1 \[\left( -\frac{ \sqrt{2} }{ 2 } \right)^2 +\left( -\frac{ \sqrt{2} }{ 2 } \right)^2 = \frac{ 2 }{ 4 }+\frac{ 2 }{ 4 }=\frac{ 4 }{ 4 }=1\]

OpenStudy (anonymous):

got it?